OpenStudy (anonymous):

Solve 9x^2 - 100 = 0 I have no idea how to solve this, and I would greatly appreciate if someone could show me in steps

jimthompson5910 (jim_thompson5910):

First factor the left side using the difference of squares rule 9x^2 - 100 = 0 (3x)^2 - (10)^2 = 0 (3x-10)(3x+10) = 0 what's next?

OpenStudy (anonymous):

find the gcf? I'm sorry I have no idea

jimthompson5910 (jim_thompson5910):

the next step would be to use the zero product property

jimthompson5910 (jim_thompson5910):

that's the idea that if A*B = 0 then A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

so if (3x-10)(3x+10) = 0 then 3x-10 = 0 or 3x+10 = 0

jimthompson5910 (jim_thompson5910):

now you would solve each individual equation for x

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

what do you get

OpenStudy (anonymous):

Sorry, I'm still figuring it out

jimthompson5910 (jim_thompson5910):

that's fine, take your time

OpenStudy (anonymous):

I think I have the final answer would it be -10/3 and 10/3?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

x = 10/3 or x = -10/3

OpenStudy (anonymous):

Thank you, very much

OpenStudy (anonymous):

Thank you for the medal too

jimthompson5910 (jim_thompson5910):

you're welcome